MsBrains (ms-brains):

Easy! Will fan and medal -3(10)÷ -3 Solve step by step The answer is 10, but how?

OpenStudy (anonymous):

just divide

MsBrains (ms-brains):

I need step by step awsners buddy. :)

OpenStudy (anonymous):

follow the order of operations. PEMDAS so -3(10)/ -3 -3*10= -30 -30/-3=10

MsBrains (ms-brains):

thank you @kitteh♥

OpenStudy (anonymous):

No problem Darling